OpenStudy (anonymous):

The slope of the line passing through the points (-2, 6) and (4, 6) is can someone please tell me how to do this??

OpenStudy (alexandervonhumboldt2):

use slope formula (y1-y2)/(x1-x2)

OpenStudy (alexandervonhumboldt2):

substitute points: (6-6)/(-2-4)=0

OpenStudy (jango_in_dtown):

The slope is given by (y2-y1)/(x2-x1) where (x1,y1) and (x2,y2) are the given points

OpenStudy (alexandervonhumboldt2):

(y1-y2)/(x1-x2) where (x1, y1) and (x2, y2) are the points: x1=-2 x2=4 y1=6 y2=6 (6-6)/(-2-4)=0

OpenStudy (anonymous):

i understand it better now thank you!